Antacids are drugs that are used to combat hyperacidity by neutralizing the excess acid (HCl) present in the stomach. Another popular component of antacid preparations is aluminum hydroxide, Al(OH)3.

1. Write a balanced chemical equation describing the neutralization reaction between aluminum hydroxide and hydrochloric acid.
2. How many grams of HCl can be neutralized by one tablespoon of an antacid preparation containing 300 mg of Al(OH)3, assuming complete reaction.
1
Expert's answer
2020-11-20T08:04:23-0500

Al(OH)3+ 3HCl=AlCl3+3H2O

n=m/M

M(Al(OH)3)=78 g/mol

M(HCl)=36.46 g/mol

n (Al(OH)3)=0.3/78=0.004 mol

n (HCl)=n (Al(OH)3) = 3 x n (Al(OH)3) = 3 x 0.004=0.012 mol

m (HCl) = n x M = 0.012 x 36.46 = 0.42 g